# VulDetectBench: Evaluating the Deep Capability of Vulnerability Detection with Large Language Models

### v3

@@ 3 Benchmark Construction
In this part, our goal is to construct a benchmark from high-quality datasets of code vulnerabilities. As shown in Table 1. We collect real-world datasets from nine popular open source projects and a synthetic dataset from the National Institute of Standards and Technology Software Assurance Reference Dataset (NIST SARD), which provides rich annotated data on code vulnerabilities. Each entry in the dataset includes a detailed classification, description and key information about the code associated with the formation and triggering of the vulnerability, facilitating the design of various tasks. We also include two open source datasets: Big-Vul[21] and Devign[22]. More detailed information about dataset can be found in Appendix.
@@ 4.2 Overall Results
Overall, we observe that existing LLMs excel in the binary classification task of detecting vulnerabilities in programs (Task 1), achieving best accuracy above 85% for inputs up to 4K tokens. This performance is consistent across both open-source and proprietary commercial models. However, their performance declines in Task 2, which involves identifying specific CWE vulnerability types. Furthermore, all models significantly underperform in tasks that require identifying key data objects and function calls associated with vulnerabilities, as well as in accurately pinpointing the root causes and trigger points of these vulnerabilities.
@@ 4.3 In-depth Analysis
Vulnerability analysis of programs using LLM has the intuitive advantage of being less dependent on the integrity of the program. Traditional static program analysis tools such as Infer require compilation of the program, which transforms program syntax trees (ASTs), program control flow graphs (CFGs), and other forms of program representation. CodeQL[11] requires the use of manually constructed databases to perform vulnerability matching. This does not allow for rapid vulnerability analysis, and we choose a commonly used tool flawfinder[12] that can directly analyze portions of a program’s code for comparative experiments. This tool can complete five tasks in VulDetectBench, but none of them give correct results and produce a large number of false positives. More detailed descriptions are provided in the Appendix.
@@ 5 Conclusion, Limitations and Societal impacts
Conclusions We have developed a new benchmark for evaluating LLMs on different tasks, using high quality real data and long program files as input. Our design of tasks of varying difficulty shows that current LLMs, both open source and proprietary, perform well on basic vulnerability detection tasks and can serve as auxiliary tools for vulnerability analysis. For example, these models are effective at detecting the presence of vulnerabilities and inferring the CWE types of the vulnerabilities. However, for more specialized tasks such as identifying key data objects and determining the root cause of vulnerabilities, their performance is still insufficient to replace traditional tools or provide effective analytical support.
@@ A.2 Details of Model Selection
Most LLMs can now process longer context lengths, but due to dataset quality and inference costs, we limit VulDetectBench’s input context to under 4K tokens with each case in dataset including only one vulnerability.Longer contexts risk including multiple vulnerabilities in a single sample.
@@ A.4 Comparision with Traditional Tools
A.5 Deep Understanding the Capability of Vulnerability Analysis
In this section,In this section, we provide detail data about the analysis of models from the same family but of varying sizeson their performance across the aforementioned five tasks.
@@ A.5 Deep Understanding the Capability of Vulnerability Analysis
A.6 Prompts of Tasks

### v4

@@ 2.2 Deep Learning in Vulnerability Detection
The work [21] proposes a framework for testing the capabilities of a given LLM as a security assistant across eight distinct dimensions. It includes the design of a multi-classification task (yes/no/n/a) for evaluating accuracy scores. Additionally, the study introduces more evaluation metrics related to the robustness of performing this task. However, it is important to note that the "root cause" identified in their work is not the specific program statement or expressions but rather the textual reasoning response generated by the LLMs. When a program contains hundreds of lines of code, such textual reasoning responses cannot precisely pinpoint the location of the vulnerability, which would be more valuable for security analysts.
Overall, the works [22] [23] [24] focus on evaluating large model vulnerability analysis capabilities. There are common problems at that time, firstly, the number of models issmall, and the evaluated models only focused on GPT4, GPT3.5 and a small number of other models such as CodeBert. Secondly, the metrics are relatively simple, and only evaluate whether the large model LLM could perform vulnerability detection on functions or code fragments. Third, there is no detailed discussion and design of the review dataset, e.g., long code, data labelling quality.
@@ 3 Benchmark Construction
In this part, our goal is to construct a benchmark from high-quality datasets of code vulnerabilities. As shown in Table 1. We collect real-world datasets from nine popular open source projects and a batch of datasets from the National Institute of Standards and Technology Software Assurance Reference Dataset (NIST SARD), which provides rich annotated data on real projects’s vulnerabilities. Each entry in the dataset includes a detailed classification, description and key information about the code associated with the formation and triggering of the vulnerability, facilitating the design of various tasks. We also include two open source datasets: Big-Vul[25] and Devign[26].
In constructing the benchmark, we prioritize data quality, referencing the work [20] to ensure the reliability of our review findings, which is not the focus in other datasets. The definition of "Accuracy" as used in[20] and in Table 2 refers to the extent to which the data attributes accurately represent the true value of the intended attribute of a concept or event. Both the Big-Vul and Devign datasets rely on vulnerability-related information scraped from GitHub commits or links to vulnerability patches for data labeling. And, after a detailed investigation, the work[20] find that the labels in these two datasets are not 100% accurate. Moreover, the Big-Vul and Devign datasets only provide labels for the existence and CWE type of vulnerability, without offering details on the root cause or trigger points. Consequently, we only selecte parts of these datasets for constructing Task 1 and Task 2 in our benchmark. Given the presence of false positives in these datasets, we exclusively included data labeled as "no vulnerability." The remaining datasets used in our benchmark are sourced from real projects, where the root cause or trigger point has been clearly annotated, ensuring 100% correct labeling. More detailed information about the datasets can be found in the Appendix A.1.
@@ 5 Vulnerability Trigger Point Location 100 38 Same Above 1011 - 3363
Using our benchmark, we feed the full code into LLMs. For task 1 and task 2, the Min-max tokens of data are 45-4051 and 132-4121, because we sample some data from Devign and Big-vul which contains many short functions. For task3, task4 and task5, all the data from real projects like Wireshark, Apache, and Openssl. The data in these datasets are very long, ranging from 1050-3300 as shown in Table2.
@@ 4.2 Overall Results
Overall, we observe that existing LLMs excel in the binary classification task of detecting vulnerabilities in programs (Task 1), achieving best accuracy above 85% for inputs up to 4K tokens. This performance is consistent across both open-source and proprietary commercial models. However, their performance declines in Task 2, which involves identifying specific CWE vulnerability types. Furthermore, all models significantly underperform in tasks that require identifying key data objects and function calls associated with vulnerabilities, as well as in accurately pinpointing the root causes and trigger points of these vulnerabilities. We also fintune the LLM to check if there are any boost of performance in Appendix A.6
@@ 4.3 In-depth Analysis
Performance on specific type of vulnerability
We conduct the experiment by testing the scores for different model models against different CWE types of vulnerabilities in task4 and task5 (in this experimental setup, as long as some lines of the answer intersected with the ground truth, the answer was judged to be primed to locate the vulnerability, which was scored as a 1, and the rest were scored as a 0, and the percentage was calculated). As shown in Table 4.
Table 4: Performance comparison of different models on Task 4 (Root Cause Location) and Task 5 (Trigger Point Location) for specific CWE types. The values represent the Unified Recall Score (URS) for each task, formatted as Task4|Task5.
CWE for Model (TASK4|TASK5) GPT4 Gemini-pro Deepseek ChatGLM3 QWEN(7B)
CWE-78 OS Command Injection 0.83|0.83 0.17|0.33 0.33|0.17 0.00|0.00 0.00|1.00
CWE-88 Argument Injection 0.71|0.14 0.71|0.14 0.29|0.57 0.00|0.29 0.57|0.57
CWE-191 Integer Underflow 0.00|0.14 0.00|0.57 0.00|0.57 0.00|0.14 0.29|0.86
CWE-89 SQL Injection 0.83|0.00 0.33|0.33 0.00|0.33 0.00|0.17 0.00|0.50
CWE-476 NULL Pointer Dereference 0.20|0.25 0.40|0.10 0.30|0.10 0.05|0.00 0.60|0.20
CWE-775 Missing Release of File Descriptor 0.60|0.80 1.00|0.20 1.00|0.20 0.00|0.00 0.80|0.60
CWE-682 Incorrect Calculation 0.00|0.00 0.25|0.50 0.25|0.00 0.00|0.00 0.00|0.25
CWE-120 Classic Buffer Overflow 1.00|0.75 0.50|1.00 0.00|0.25 0.25|0.25 0.00|0.75
CWE-190 Integer Overflow or Wraparound 0.00|0.00 0.33|0.00 0.00|0.33 0.00|0.00 0.00|0.68
We also conduct the following experiment: for each CWE type in TASK2 (top 10 in number of 500 test cases, num is the specific number) the model corresponds to the proportion of correctly classified, as shown in Table 5.
Table 5: Performance comparison of different models on Task 2 (Vulnerability CWE Type Inference) for the top 10 most frequent CWE types. The table shows the proportion of correctly classified instances for each CWE type across various models. The ’Num’ column indicates the number of test cases for each CWE type out of 500 total test cases.
CWE TYPE Num GPT4 Gemini-pro Deepseek ChatGLM3 Qwen(7B)
CWE-78 OS Command Injection 75 1.00 0.77 0.66 0.24 0.77
CWE-90 LDAP Injection 74 0.99 0.78 0.63 0.14 0.79
CWE-23 Relative Path Traversal 72 1.00 0.77 0.42 0.09 0.80
CWE-36 Absolute Path Traversal 68 1.00 0.91 0.26 0.13 0.88
CWE-114 Process Control 43 1.00 0.48 0.28 0.22 0.16
CWE-15 External Control of System or Configuration Setting 37 1.00 0.97 0.43 0.07 0.70
CWE-476 NULL Pointer Dereference 26 0.62 0.44 0.14 0.04 0.33
CWE-89 SQL Injection 8 1.00 0.88 0.13 0.00 0.25
CWE-191 Integer Underflow (Wrap or Wraparound) 8 0.50 0.50 0.00 0.00 0.31
CWE-88 Argument Injection 7 0.93 0.79 0.14 0.14 0.43
The results of these two experiments indicate that LLMs exhibit varying abilities in analyzing different specific types of vulnerabilities. In Task 2, the closed-source model demonstrates superior overall performance, though the open-source model is also capable of handling certain tasks, particularly for vulnerabilities with more distinctive features, such as CWE-78 and CWE-23. LLMs show competence in identifying CWE types in these cases. However, for Tasks 4 and 5, only GPT-4 is able to provide relatively valuable vulnerability localization results for specific tasks. For vulnerability types like CWE-191, which require procedural analysis and possess more hidden characteristics, LLMs are not yet capable of delivering valuable information.
@@
Vulnerability analysis of programs using LLM has the intuitive advantage of being less dependent on the integrity of the program. Traditional static program analysis tools such as Infer require compilation of the program, which transforms program syntax trees (ASTs), program control flow graphs (CFGs), and other forms of program representation. CodeQL[11] requires the use of manually constructed databases to perform vulnerability matching. This does not allow for rapid vulnerability analysis, and we choose a commonly used tool flawfinder[12] that can directly analyze portions of a program’s code for comparative experiments. This tool can complete five tasks in VulDetectBench, but none of them give correct results and produce a large number of false positives. More detailed descriptions are provided in the Appendix A.4.
@@ 5 Conclusion, Limitations and Societal Impacts
Conclusions Our study introduces a new benchmark for evaluating LLMs on a range of tasks using high-quality real data and long program files as input. By focusing on tasks of increasing complexity, we have gained valuable insights into the capability boundaries of both open-source and proprietary LLMs. While these models perform well on basic vulnerability detection tasks—effectively detecting the presence of vulnerabilities and inferring CWE types—they fall short in more specialized tasks, such as identifying key data objects and accurately determining the root cause of vulnerabilities. Our research emphasizes the need for precision in root cause identification, particularly when dealing with extensive codebases, where pinpointing exact vulnerability locations is essential. Additionally, our evaluation methodology directly assesses the LLMs’ ability to locate root causes, trigger points, or key data objects, offering a more granular and precise assessment of their performance. Although LLMs can serve as useful auxiliary tools for vulnerability analysis, their current capabilities are insufficient to replace traditional tools or provide comprehensive analytical support in more complex scenarios.
@@ A.2 Details of Model Selection
Most LLMs can now process longer context lengths, but due to dataset quality and inference costs, we limit VulDetectBench’s input context to under 4K tokens with each case in dataset including only one vulnerability. Longer contexts risk including multiple vulnerabilities in a single sample. However, as models evolve and support longer window size, we have designed a dataset with longer code length for evaluation including 8k, 16k, 32k, etc, and will update it to the repository.
@@ A.4 Comparision with Traditional Tools
For static testing tools, like Flawfinder presented, it just treats the program as text and uses fixed patterns to detect and give a report of vulnerability. And we can see it makes lots of false positives and hardly provides useful information for security analysts. So, to conduct a more precise program analysis, like inter-procedure program analysis and context-sensitive analysis and so on, the program need to be transfered into AST(abstract syntax trees) or some IR representaions(like LLVM IR) like CodeQL [11] and Cppcheck[13] . That requires the program are complete and can be successfully compiled.
A.5 Deep Understanding of the Capability of Vulnerability Analysis
In this section,we provide detailed data about the analysis of models from the same family but of varying sizeson their performance across the aforementioned five tasks.
@@ A.5 Deep Understanding of the Capability of Vulnerability Analysis
A.6 Finetuing the LLM
At the same time, we have carried out finetuning experiments, which mainly focused on tasks 3, 4 and 5, which had poor performance of the fundamental model. Task 3 was the intermediate analysis process of tasks 4 and 5, so we could directly fine-tune tasks 4 and 5 without considering task 3. Through the analysis of the experiment of fine-tuning Meta-Lema-3-8B-Instruct, we found that the fundamental model basically did not show any ability for the two tasks. The model used in this evaluation is the Meta-Llama-3-8B-Instruct, which is designed for the CAUSAL_LM task type. The rank is set to 8. The model is fine-tuned using a LoRA (Low-Rank Adaptation) with an alpha value of 32 and a dropout rate of 0.1. The learning rate applied during fine-tuning is 2e-4, and the model is trained over 5 epochs.
Table 11: Performance of Fine-tuned Models on TASK4 and TASK5
Model TASK4 TASK5
Meta-Llama-3-8B-Instruct 01.10 00.53
Meta-Llama-3-8B-Instruct (Fine-tuned) 39.00 52.07
After fine-tuning, the recall rate increased to 39.00% for task4, and 52.07% for task5. we can see that there is a definite improvement. However, due to strict evaluation requirements, the effect is only improved to a certain extent under the current metric. We conducted an in-depth analysis of the model’s responses and found that fine-tuning significantly enhanced its ability to perceive vulnerability locations. However, pinpointing a specific line remains a challenging task, and we are continuing to work on it.
A.7 Chain of Thought and Few-shot Testing
Furthermore, we conducted additional experiments to investigate the model’s performance on vulnerability detection tasks under different prompting technique settings. We used GPT-4 and tested it on tasks 3, 4, and 5 using both Chain of Thought (CoT) and Few-shot prompting techniques. The results are as follows.
Table 12: Experimental results for vulnerability detection tasks using CoT.
Method TASK3 TASK4 TASK5
Zero-shot 20.21|16.07 24.26|27.07 13.00|17.85
CoT 45.54|39.85 1.25|16.77 1.64|22.88
Few-shot (2-shot) 10.34|13.41 9.64|9.64 15.37|15.37
Experimental results show that CoT improves vulnerability analysis, while few-shot prompting may negatively impact localization. We attribute this to the following reasons:
  1. 1.
    Vulnerability diversity makes few-shot examples potentially misleading.
  2. 2.
    CoT, designed based on expert analysis methods, aligns with standard vulnerability assessment paradigms, improving performance in tasks 3 and 5.
A.8 Prompt Design